Ok I need steps for the question " a store allows customers to buy 3 iteams of the same price and get 2 more of the same iteams free. What percent savings does this represent.
The answer is 40 but i need steps im only 12 so make the steps easy. thanks :p

Found 2 solutions by scott8148, josmiceli:
Answer by scott8148(6628)

putting in numbers may make it easier

3 items at $1.00 each, plus 2 free is 5 items for $3.00

5 items for $3.00 is $.60 each ($3.00 / 5)

the $.60 price is 60% of the $1.00 price [($.60 / $1.00) * 100%]

100% (full price) minus 60% (sale price) equals 40% savings

Answer by josmiceli(19441)

First find the savings.
This is ( amount you would normally pay ) - ( amount you actually pay )
--------------------------
Then to find the percent savings,
% savings = ( savings ) / ( amount you you normally pay )
and multiply this fraction by 100
-------------------------
Let +c+ = normal cost of 5 items
savings = c - (3/5)*c
( note that (3/5)*c means customer only paid for 3 of the 5 items )
----------------
c - (3/5)*c = (2/5)*c
% savings = [( 2/5)*c / c ] x 100
The c's cancel
% savings = (2/5)x 100
% savings = 40
---------------------
If this is a little hard to follow, the most
important part is really
% savings = ( savings ) / ( amount you you normally pay )
then multiply by 100